# Flamelet Connection to Turbulence Kinetic Energy Dissipation Rate

### v1

@@ 7 Scaling of Velocity Gradients
The positive coefficient needs a value $C_{ke}<C_{vd}$ in order for the counterflow to exist. A given value of $\epsilon$ will produce one acceptable $\omega$ value and one acceptable $S^{*}$ value. Thereby, we expect one acceptable $SDR_{max}$ value. However, that value cannot be captured by knowing $S^{*}$ alone.
It is found that $\epsilon$ is the meaningful tracking variable. From the knowledge of $\epsilon$ and fluid properties, with a choice for the parameter $S_{1}$, the values of $S^{*}$, dimensional $\omega,$ the pressure Laplacian, and the viscous dissipation follow. Specifically,
@@
We must be sure that, with any application of the flamelet model for Kolmogorov or near-Kolmogorov scales, $(1/2)C_{ke}<(3/4)C_{vd}$ is at least supportable in a statistical sense. No counterflow-based flamelet model can be justified with a positive pressure Laplacian.
Using the resolved-scale non-dimensionalization scheme discussed in Appendix B, it can be shown that
$\frac{S*}{\beta_{\infty}}$ $=$ $\frac{1}{2\beta_{\infty}}\Bigg{(}\frac{C_{vd}\;\epsilon}{\nu[S_{1}^{2}+1-S_{1}]}\Bigg{)}^{1/2}=\frac{1}{2}\Bigg{(}\frac{C_{vd}\;Re\;\epsilon_{nd}}{S_{1}^{2}+1-S_{1}}\Bigg{)}^{1/2}\;\;;\;\;$
@@
Plots of the flammability curve for the rotational flamelet, with curves of nondimensional $T_{max},SDR_{max}$, or integrated burning rate versus $\sqrt{\epsilon/\nu_{\infty}}$, are informative, e.g., plots of the Hellwig et al. (2023a) flamelet results (integrated burning rate, $T_{max}$, and $SDR_{max}$) versus $\epsilon$ at a fixed $S_{1}$.
$\epsilon$ should not be considered a “progress variable” since its value can increase or decrease in any spacial direction or in time. We should think in terms of a “tracking variable” rather than a progress variable.
@@ 8 Flamelet Connection with $\epsilon$
In this section, results from five flamelet models for $\mathrm{H_{2}/N_{2}-O_{2}}$ combustion are presented. These models are generated for an oxidizer stream consisting of pure $\mathrm{O_{2}}$ and a fuel stream consisting of an equimolar mixture of $\mathrm{H_{2}}$ and $\mathrm{N_{2}}$. Both streams enter at 300 K, with a background pressure of 10 atm. Gas-phase kinetics, thermal properties, and transport properties are calculated using a nine-species reduction of Version 1.0 of the Foundational Fuel Chemistry Model reaction mechanism (FFCM-1) developed and tested in Smith et al. (2016); Tao et al. (2018). Nitrogen is treated as an inert species.
Figure 2: Comparison of three flamelet models for $\mathrm{H_{2}/N_{2}-O_{2}}$ diffusion flames with $T_{max}$ as a function of stoichiometric SDR value ($\chi_{st}$): red, Progress Variable result from FlameMaster; black, Rotational Flamelet with Fickian diffusion at Le=1 and zero vorticity; blue, Rotational Flamelet with Fickian diffusion at Le=1 and vorticity.
The first set of results pertains to the classical flamelet model, where key quantities of interest—such as temperature, species mass fractions, and heat release—are determined by solving the classical flamelet equations in mixture fraction space (Peters (1984, 2000); Pierce & Moin (2004)). In this context, the mixture fraction is denoted by $Z$. Unlike the Rotational Flamelet model, the momentum equation is not explicitly solved in the classical flamelet model. Instead, its effects are introduced through an assumed form of the scalar dissipation rate, specifically $\chi(Z)=\frac{2S^{*}}{\pi}\exp{(2\mathrm{erfc}^{-1}(2Z)^{2})}$ as specified in Peters (2000). Note that by not explicitly solving the momentum equation, this formulation does not address the effects of vorticity on the flame. The results for the classical flamelet model are obtained using the FlameMaster code developed by Pitsch (2022).
Figure 3: Comparison of five flamelet models for $\mathrm{H_{2}/N_{2}-O_{2}}$ diffusion flames with $T_{max}$ as a function of applied compressive strain rate $S^{*}$: red, Progress Variable result from FlameMaster; black, Rotational Flamelet with Fickian diffusion at Le=1 and zero vorticity; blue, Rotational Flamelet with Fickian diffusion at Le=1 and vorticity; dashed black, Rotational Flamelet with preferential diffusion and zero vorticity; dashed blue, Rotational Flamelet with preferential diffusion and vorticity.
The remaining four sets of results correspond to Rotational Flamelet models. These include Rotational Flamelet models with and without vorticity using Fickian diffusion at unity Lewis number, as well as models with and without vorticity using preferential mass diffusion. These rotational flamelet models were generated using an in-house MATLAB code developed in Hellwig et al. (2023a, b). To ensure optimal comparability between the classical and Rotational Flamelet models, equal transverse strain rates ($S_{1}$ = $S_{2}$) are employed in the rotational model computations. That is, $S_{1}$ is set to 1/2 in the equations presented above. This value provides the closest approximation to the classical model’s two-dimensional axisymmetric counterflow configuration.
The results focus on three main parameters: maximum flamelet temperature, integrated burning rate, and stoichiometric flamelet scalar dissipation rate. These parameters are examined as functions of both inflow strain rate $S^{*}$ and $\epsilon$. The latter serves to highlight the connection between the flamelet and the resolved scale. Note that in the previous sections, unlike some other works, $\omega$ is dimensional with identical units as $S^{*}$. However, in this section, $\omega$ is non-dimensionalized by $S^{*}$, following the approaches used by Sirignano (2022c, a, b) and Hellwig et al. (2023a, b).
Figure 4: Comparison of five flamelet models for $\mathrm{H_{2}/N_{2}-O_{2}}$ diffusion flames with $T_{max}$ as a function of $\epsilon/\nu_{\infty}$: red, classical result from FlameMaster; black, Rotational Flamelet with Fickian diffusion at Le=1 and zero vorticity; blue, Rotational Flamelet with Fickian diffusion at Le=1 and vorticity; dashed black, Rotational Flamelet with preferential diffusion and zero vorticity; dashed blue, Rotational Flamelet with preferential diffusion and vorticity.
@@
Figure 2 presents the maximum flamelet temperature as a function of the stoichiometric scalar dissipation rate, commonly known as S-shaped curves. This figure serves as a common reference for readers familiar with flamelet modeling based on the classical mixture fraction formulation. Only three curves are shown, as the scalar dissipation rate is not applicable in cases of preferential diffusion. Differences between the classical and rotational models are evident and anticipated, reflecting the variations in their formulations. The rotational models demonstrate marginally extended flammability limits for both vorticity values. While this presentation, which parametrizes the flame state using scalar dissipation rate, may suggest that vorticity has a negligible impact, it will be shown later that relying on $\chi$ to describe the flame state is misleading and obscures the significant implications of accounting for vorticity.
Figure 5: Comparison of three flamelet models for $\mathrm{H_{2}/N_{2}-O_{2}}$ diffusion flames with $\chi_{st}$ as a function of $S^{*}$: red, classical result from FlameMaster; blue, Rotational Flamelet with Fickian diffusion at Le=1 and zero vorticity; black, Rotational Flamelet with Fickian diffusion at Le=1 and vorticity.
@@
Figure 3 displays the same S-shaped curves, but now parameterized by the inflow strain rate $S^{*}$. For the classical model, the S-shaped curve plotted against strain rate remains qualitatively similar to that in Figure 2. This is because $S^{*}$ is related to $\chi_{st}$ by a constant through assumed functional form of $\chi(Z)$. However, in the case of the rotational models—where $\chi(Z)$ is not assumed but instead determined by solving the momentum equation—the effect of vorticity on the flamelet becomes more pronounced. The zero-vorticity case closely resembles the classical flamelet model, which is expected, as the absence of vorticity aligns with the classical formulation where vorticity is not considered. However, when vorticity is included, a significant increase in the extinction inflow strain rate is observed.
Additionally, two more curves are plotted for Rotational Flamelet models with preferential diffusion, corresponding to dimensionless vorticity values of zero and one. Once again, the presence of vorticity extends the flammability limit. Moreover, the use of preferential diffusion results in distinct differences, including increased maximum temperature and broader flammability limits compared to the Fickian diffusion models.
Figure 6: Comparison of three flamelet models for $\mathrm{H_{2}/N_{2}-O_{2}}$ diffusion flames with $\chi_{st}$ as a function of $\epsilon/\nu_{\infty}$: red, classical result from FlameMaster; black, Rotational Flamelet with Fickian diffusion at Le=1 and zero vorticity; blue, Rotational Flamelet with Fickian diffusion at Le=1 and vorticity.
@@
In Figure 4, the maximum temperature is parameterized by $\epsilon/\nu_{\infty}$. To achieve this, $S^{*}$ is related to $\epsilon/\nu$ using the first expression from Equation 7. Although the value of $C_{vd}$ is not known at this time, a value of 1 has been selected for demonstration purposes. With $\mathrm{S_{1}}$=1/2, the inflow strain rate becomes $S^{*}=\sqrt{\epsilon/(3\nu)}$. This illustrates how $\epsilon$, which is determined from the resolved scale, can serve as a tracking effectively connecting the flamelet to resolved scale computations.
Figure 7: Comparison of three flamelet models for $\mathrm{H_{2}/N_{2}-O_{2}}$ diffusion flames with $\chi_{st}$ as a function of $\sqrt{\epsilon/\nu_{\infty}}$: red, classical result from FlameMaster; black, Rotational Flamelet with Fickian diffusion at Le=1 and zero vorticity; blue, Rotational Flamelet with Fickian diffusion at Le=1 and vorticity.
Walsh et al. (2024) employed a Flamelet Progress Variable (FPV) approach coupled with the RANS equations to study methane-vitiated air combustion in highly accelerating mixing layers. In the FPV approach, $\chi_{st}$ is monotonically mapped to a progress variable that is defined at the sub-scale but tracked at the resolved scale using a constructed transport equation. However, as shown in Figure 4, it is conceivable to replace the progress variable with $\epsilon$, which is directly available from the resolved-scale RANS equations.
Figure 8: Comparison of three flamelet models for $\mathrm{H_{2}/N_{2}-O_{2}}$ diffusion flames with integrated burning rate as a function of $S^{*}$: red, classical result from FlameMaster; black, Rotational Flamelet with Fickian diffusion at Le=1 and zero vorticity; blue, Rotational Flamelet with Fickian diffusion at Le=1 and vorticity.
@@
Figures 5 and 6 plot stoichiometric scalar dissipation rates $\chi_{st}$ for the three Fickian models against $S^{*}$ and $\epsilon/\nu_{\infty}$, respectively. While $\chi_{st}$ has been chosen for consistency with the classical models, there is no inherent reason why $\chi_{max}$ could not be used instead. The $\chi_{st}$ used in the classical model, represented by the red line, is bijective and grows linearly with $S^{*}$, as a result of the assumed functional form of $\chi(Z,\omega)$. As previously discussed, in the Rotational Flamelet model, scalar dissipation is not assumed but determined by solving the momentum equation. The differences between the blue and black curves clearly demonstrate how scalar dissipation rate becomes a function of vorticity, i.e., $\chi(Z,\omega)$. Furthermore, contrary to the classical model, stable and unstable solutions have distinct $\chi_{st}$ values.
@@
This dependence of $\chi$ on $\omega$ has significant implications. When viewing these figures alongside Figure 2, it becomes evident that for a given value of $\epsilon$ determined from the resolved scale, which imposes a corresponding $S^{*}$ value, different vorticity values result in two distinct $\chi_{st}$, values, leading to two different flamelet solutions along the S-shaped curve. These solutions exhibit different maximum temperatures and integrated burning rates, regardless of the fact that the S-shaped curves for different vorticity values collapse onto each other when parameterized by $\chi_{st}$. This clearly shows that a flamelet solution cannot be determined based on $S^{*}$ alone.
@@
Figure 7 shows the same $\chi_{st}$ curves plotted against the square root of $\epsilon/\nu_{\infty}$. $\epsilon$ is the rate of dissipation of k, the turbulence kinetic energy per unit mass. It follows that the magnitude of velocity fluctuation decays at a rate proportional to $\sqrt{\epsilon}$. If we assume that a scalar fluctuation decreases proportionately to velocity and $\chi$ represents a rate of dissipation for a scalar fluctuation, the proportionality of $\chi$ and $\sqrt{\epsilon}$ are observed and expected.
Figure 9: Comparison of three flamelet models for $\mathrm{H_{2}/N_{2}-O_{2}}$ diffusion flames with integrated burning rate as a function of $\epsilon/\nu_{\infty}$: red, classical result from FlameMaster; black, Rotational Flamelet with Fickian diffusion at Le=1 and zero vorticity; blue, Rotational Flamelet with Fickian diffusion at Le=1 and vorticity.
@@
More relevant than flame temperature is the heat release, which will in practice be the parameter extracted from the flamelet model and introduced into the resolved-scale computations. Figures 8 and 9 depict the integrated burning rate for the three Fickian models as functions of $S^{*}$ and $\epsilon/\nu_{\infty}$, respectively. Here, the integrated burning rate is calculated as the heat release rate integrated over $Z$. All three models show that the integrated burning rate increases in a nearly linear manner with $S^{*}$, indicating that it grows as the scale decreases. As expected, the Rotational Flamelet model with zero vorticity most closely resembles the classical model, with differences in integrated burning rates due to distinctions in the underlying formulations of the two models. In the presence of vorticity (blue curve), the integrated burning rate shows a significant decrease compared to the zero-vorticity case.
Hellwig et al. (2023b) observed similar behaviour for integrated burning rate of $\mathrm{H_{2}O}$ plotted against $S^{*}$, showing a decrease in burning rates with increasing vorticity magnitude. Additionally, they found that when this parameter is plotted against $\chi_{st}$, all cases with different vorticity magnitude collapse onto a single curve. However, as was show in Figures 2 and 5, it is the combination of knowledge of both $S^{*}$ and $\omega$ (or simply $\epsilon$) that determines an appropriate value of $\chi_{st}$, leading to different points along these collapsed curves.
Figure 10: Comparison of three flamelet models for $\mathrm{H_{2}/N_{2}-O_{2}}$ diffusion flames with $SDR$ as a function of mixture fraction $Z$: red, classical from FlameMaster; black, Rotational Flamelet with Fickian diffusion at Le=1 and zero vorticity; blue, Rotational Flamelet with Fickian diffusion at Le=1 and vorticity.
Figure 11: Comparison of three flamelet models for $JP5$-Air diffusion flames with $SDR$ as a function of mixture fraction $Z$: red, classical result from FlameMaster; black, Rotational Flamelet with Fickian diffusion at Le=1 and zero vorticity; blue, Rotational Flamelet with Fickian diffusion at Le=1 and vorticity.
@@
Figures 10 and 11 are examples of scalar dissipation rate $\chi$, versus mixture fraction $Z$, for $\mathrm{H_{2}},\mathrm{O_{2}},\mathrm{N_{2}}$ combustion and JP-5, air combustion, respectively. The blue and black curves illustrate the difference in scalar dissipation rate created by solving the momentum equations with vorticity in addition to the energy and species continuity equations. Clearly, vorticity has an effect on scalar dissipation rate and will thus alter the solution of the flamelet equations.
@@
Furthermore, the peak of the scalar dissipation rate curve shows bias towards the stream with higher thermal diffusivity. This is seen from the $\mathrm{H_{2}},\mathrm{O_{2}},\mathrm{N_{2}}$ peak sloping toward the fuel stream while the JP-5 peaks slopes away from the fuel stream. For Le=1, the mass diffusivity value is set to the thermal diffusivity value which is the product of density and specific heat in the denominator. Both terms in that denominator product minimize on the fuel-rich (fuel-lean) side for $\mathrm{H_{2}}/\mathrm{N_{2}-O_{2}}$ (JP5-Air), causing diffusivity to maximize on that side. This results in $dZ/dy$, its square, and the coefficient $D$ maximizing and causing $\chi$ to maximize on that preferred side. Let us, for the sake only of argument and display here, consider that, in the vicinity of maximum SDR, the velocity $v=-S^{*}y$ and $D$ is approximately constant over a neighborhood. Treatment of the second-order ODE for $Z(y)$ as a first-order ODE for $dZ/dy$ yields a local behavior where $dZ/dy\propto\exp{[-S^{*}y^{2}/(2D)]}$ and $\chi\propto D\exp{[-S^{*}y^{2}/D]}$. The larger $D$ in the denominator of the exponent causes a slower decline from the maximum value and thereby a larger value of the exponential term as well as its coefficient. This difference in thermal diffusivity for unitary Lewis number is due to the relatively high specific heat for JP-5 and $\mathrm{O_{2}}$ relative to their respective opposing streams.
Since the scalar dissipation rate is proportional to strain rate, we should expect $SDR_{max}/S^{*}\propto\epsilon^{1/2}$ and scalar gradients should behave as $\epsilon^{1/4}$.

### v2

@@ 6 Scaling of Velocity Gradients
It is required that $C_{ke}<C_{vd}$ in order for the counterflow to exist with pressure peaking at a stagnation point. Furthermore, we must have a positive enstrophy in Equation (26). Thus, the complete requirement is $C_{vd}/2<C_{ke}<C_{vd}$. A given value of $\epsilon$ will produce one acceptable $\omega$ value and one acceptable $S^{*}$ value. Thereby, we expect one acceptable SDRmax value. However, that value cannot be captured by knowing $S^{*}$ alone.
Thus, $\epsilon$ is the meaningful tracking variable. From the knowledge of $\epsilon$ and fluid properties, with a choice for the parameter $S_{1}$, the values of $S^{*}$, dimensional $\omega,$ the pressure Laplacian, and the viscous dissipation follow. Specifically,
@@
We must be sure that, with any application of the flamelet model for Kolmogorov or near-Kolmogorov scales, $C_{vd}/2<C_{ke}<C_{vd}$ is at least supportable in a statistical sense. No counterflow-based flamelet model can be justified with a positive pressure Laplacian.
It can be shown that
@@
$\epsilon$ should not be considered a “progress variable” since its value can increase or decrease in any spacial direction or in time. We should think in terms of a “coupling variable” rather than a progress variable.
@@ 7 Flamelet Connection with $\epsilon$
In this section, the mapping of flamelet solutions to the turbulent kinetic energy dissipation rate $\epsilon$ is presented for two fuel–oxidizer combinations: $\mathrm{H_{2}/N_{2}-O_{2}}$ and JP-5–air. For the $\mathrm{H_{2}/N_{2}-O_{2}}$ configuration, the oxidizer stream consists of pure $\mathrm{O_{2}}$, while the fuel stream is an equimolar mixture of $\mathrm{H_{2}}$ and $\mathrm{N_{2}}$. Both streams enter at 300 K, and the background pressure is set to 10 atm. Gas-phase thermochemistry and transport properties for the $\mathrm{H_{2}-O_{2}}$ case are computed using a nine-species skeletal reduction of Version 1.0 of the Foundational Fuel Chemistry Model (FFCM-1) [Smith2016, TAO201818], where nitrogen is treated as inert. For the JP-5–air configuration at 20 atm, the HyChem A3 skeletal reaction mechanism, which includes 40 chemical species, is employed [wang_physics-based_2018, xu_physics-based_2018].
Three flamelet models are evaluated: a steady, classical mixture-fraction-based flamelet model, and two variations of the Rotational Flamelet Model (RFM), including models with and without vorticity. All models assume Fickian diffusion at unity Lewis number. Although unsteady flamelet formulations are more suitable for capturing transient phenomena such as ignition and extinction [sirignano_unsteady_2024], the present analysis focuses on steady-state formulations to the viability of using $\epsilon$ as a coupling variable. The conclusions drawn here remain applicable in the unsteady context.
The classical flamelet solutions are computed in mixture fraction space using the steady laminar flamelet equations [Peters, Peters1984], with mixture fraction denoted by $Z$. In contrast to the RFM, the momentum equation is not explicitly solved in the classical model. Instead, its effects are introduced by assuming the standard scalar dissipation rate profile
$\chi(Z)=\frac{2S^{*}}{\pi}\exp\left(-2\left[\mathrm{erfc}^{-1}(2Z)\right]^{2}\right)$ (30)
This slightly differs from Peters [Peters] by a factor of the square root of the ratio of the two densities for the incoming streams, because, we use the strain rate of the incoming fuel stream rather than the oxidizer stream. As the classical formulation does not include a momentum equation, vorticity effects are inherently excluded. These solutions are generated using the FlameMaster code [flamemaster].
The remaining two solution sets correspond to Rotational Flamelet Models, generated using an in-house MATLAB solver developed in previous work [Hellwig2025a, Hellwig2025b]. To ensure meaningful comparison with the CFM, the RFM calculations employ equal transverse strain rates ($S_{1}=S_{2}$), with $S_{1}=1/2$ to replicate the two-dimensional axisymmetric counterflow configuration characteristic of the classical flamelet formulation. Naturally, the best comparison with the CFM is for the zero vorticity RFM cases.
The analysis focuses on three key output quantities: the maximum flamelet temperature, the integrated heat release rate (burning rate), and the stoichiometric scalar dissipation rate. These quantities are evaluated as functions of both the inflow strain rate $S^{*}$ and the turbulent kinetic energy dissipation rate $\epsilon$. The objective here is not to assess whether the rotational flamelet model is superior to the classical flamelet formulation. Results from the classical model are included primarily for reference, as it remains more familiar to many readers. Rather, the focus is on demonstrating how the turbulent kinetic energy dissipation rate $\epsilon$ may serve as a viable coupling variable in both modeling frameworks, and on highlighting how its role differs (particularly with respect to the treatment of vorticity) in the classical versus rotational flamelet approaches.
@@
Figure 2 presents the maximum flamelet temperature as a function of the stoichiometric scalar dissipation rate, commonly known as S-shaped curves. This figure serves as a common reference for readers familiar with flamelet modeling based on the classical mixture fraction formulation. Differences between the classical and rotational models are evident and anticipated, reflecting the variations in their formulations. The rotational models demonstrate marginally extended flammability limits for both vorticity values. While this presentation, which parametrizes the flame state using scalar dissipation rate, may suggest that vorticity has a negligible impact, it will be shown later that relying on $\chi(Z)$ to describe the flame state is misleading and obscures the significant implications of accounting for vorticity.
@@
Figure 3 displays the same S-shaped curves, but now parameterized by the inflow strain rate $S^{*}$. For the classical model, the S-shaped curve plotted against strain rate remains qualitatively similar to that in Figure 2. This occurs because $S^{*}$ is related to $\chi_{st}$ by a constant through assumed functional form of $\chi(Z)$. However, in the case of the rotational models (where $\chi(Z)$ is not assumed but instead determined by solving the momentum equation) the effect of vorticity on the flamelet becomes more pronounced. The zero-vorticity case closely resembles the classical flamelet model, which is expected, as the absence of vorticity aligns with the classical formulation where vorticity is not considered. However, when vorticity is included, a significant increase in the extinction inflow strain rate is observed.
@@
In Figure 4, the maximum temperature is parameterized by $\epsilon/\nu_{\infty}$. To achieve this, $S^{*}$ is related to $\epsilon/\nu$ using the first expression from Equation 28. Although the value of $C_{vd}$ is not known at this time, a value of 1 has been selected for demonstration purposes. With $S_{1}=1/2$, the inflow strain rate becomes $S^{*}=\sqrt{\epsilon/(3\nu)}$. This illustrates how $\epsilon$, which is determined from the resolved scale, can serve as a tracking effectively connecting the flamelet to resolved scale computations.
@@
Figures 5 and 6 plot stoichiometric scalar dissipation rates $\chi_{st}$ for the three models against $S^{*}$ and $\sqrt{\epsilon/\nu_{\infty}}$, respectively. The $\chi_{st}$ used in the classical model, represented by the red line, is bijective and grows linearly with $S^{*}$, as a result of the assumed functional form of $\chi(Z,\omega)$. As previously discussed, in the Rotational Flamelet model, scalar dissipation is not assumed but determined by solving the momentum equation. The differences between the blue and black curves clearly demonstrate how scalar dissipation rate becomes a function of vorticity, i.e., $\chi(Z,\omega)$. Furthermore, contrary to the classical model, stable and unstable solutions have distinct $\chi_{st}$ values.
@@
This dependence of $\chi(Z)$ on $\omega$ has significant implications. When viewing these figures alongside Figure 2, it becomes evident that, for a given value of $\epsilon$ determined from the resolved scale, which imposes a corresponding $S^{*}$ value, different vorticity values result in two distinct $\chi_{st}$, values, leading to two different flamelet solutions along the S-shaped curve. These solutions exhibit different maximum temperatures and integrated burning rates, regardless of the fact that the S-shaped curves for different vorticity values collapse onto each other when parameterized by $\chi_{st}$. This clearly shows that a flamelet solution cannot be determined based on $S^{*}$ alone.
@@
Figure 6 shows the same $\chi_{st}$ curves plotted against the square root of $\epsilon/\nu_{\infty}$. $\epsilon$ is the rate of dissipation of k, the turbulence kinetic energy per unit mass. It follows that the magnitude of velocity fluctuation decays at a rate proportional to $\sqrt{\epsilon}$. If we assume that a scalar fluctuation decreases proportionately to velocity and $\chi(Z)$ represents a rate of dissipation for a scalar fluctuation, the proportionality of $\chi(Z)$ and $\sqrt{\epsilon}$ are observed and expected.
@@
More relevant than flame temperature is the heat release, which will in practice be the parameter extracted from the flamelet model and introduced into the resolved-scale computations. Figures 7 and 8 depict the integrated burning rate for the three models as functions of $S^{*}$ and $\sqrt{\epsilon/\nu_{\infty}}$, respectively. Here, the integrated burning rate is calculated as the heat release rate integrated over $Z$. All three models show that the integrated burning rate increases in a nearly linear manner with $S^{*}$, indicating that it grows as the scale decreases. As expected, the Rotational Flamelet model with zero vorticity most closely resembles the classical model, with differences in integrated burning rates due to distinctions in the underlying formulations of the two models. In the presence of vorticity (blue curve), the integrated burning rate shows a significant decrease compared to the zero-vorticity case.
Figures 9, 10, and 11 show key results for the jet fuel JP-5 burning with air. The JP-5 air models are generated for an oxidizer stream consisting of pure air and a fuel stream consisting of pure JP-5. Both streams enter at 300 K, with a background pressure of 20 atm. Gas-phase kinetics, thermal properties, and transport properties are calculated using the A3 skeletal version of the HyChem Combustion reaction models [wang_physics-based_2018, xu_physics-based_2018]. Nitrogen is treated as an inert species. We see that JP-5 with its much slower chemical reaction rate than hydrogen will not sustain as high a strain rate; the flammability domain is lower by several orders of magnitude for any variable of interest $S^{*}$, $\epsilon$, or $\sqrt{\epsilon}$. On the other hand, the percentage increase of flammability limit due to vorticity is substantially greater.
@@
Figures 12 and 13 are examples of scalar dissipation rate $\chi(Z)$, versus mixture fraction $Z$, for $\mathrm{H_{2}},\mathrm{O_{2}},\mathrm{N_{2}}$ combustion and JP-5, air combustion, respectively. The blue and black curves illustrate the difference in scalar dissipation rate created by solving the momentum equations with vorticity in addition to the energy and species continuity equations. Clearly, vorticity has an effect on scalar dissipation rate and will thus alter the solution of the flamelet equations.
@@
Furthermore, the peak of the scalar dissipation rate curve shows bias towards the stream with higher thermal diffusivity. This is seen from the $\mathrm{H_{2}},\mathrm{O_{2}},\mathrm{N_{2}}$ peak sloping toward the fuel stream while the JP-5 peaks slopes away from the fuel stream. For $Le=1$, the mass diffusivity value is set to the thermal diffusivity value which has the product of density and specific heat in the denominator. Both terms in that denominator product minimize on the fuel-rich (fuel-lean) side for $\mathrm{H_{2}}/\mathrm{N_{2}-O_{2}}$ (JP-5 - Air), causing diffusivity to maximize on that side. This results in $dZ/dy$, its square, and the coefficient $D$ maximizing and causing $\chi(Z)$ to maximize on that preferred side. Let us, for the sake only of argument and display here, consider that, in the vicinity of maximum SDR, the velocity $v=-S^{*}y$ and $D$ is approximately constant over a neighborhood. Treatment of the second-order ODE for $Z(y)$ as a first-order ODE for $dZ/dy$ yields a local behavior where $dZ/dy\propto\exp{[-S^{*}y^{2}/(2D)]}$ and $\chi\propto D\exp{[-S^{*}y^{2}/D]}$. The larger $D$ in the denominator of the exponent causes a slower decline from the maximum value and thereby a larger value of the exponential term as well as its coefficient. This difference in thermal diffusivity for unitary Lewis number is due to the relatively high specific heat for JP-5 and $\mathrm{O_{2}}$ relative to their respective opposing streams.
[…]
8 Conclusions
The turbulence kinetic energy dissipation rate $\epsilon$ have been defined at the resolved scale of turbulent flow fields, including flows with turbulent combustion. $\epsilon$ is readily defined in both RANS and detached-eddy simulations. It is an excellent parameter for closure with sub-grid non-premixed flamelet models to determine the vorticity $\omega$ and the compressive normal strain rate $S^{*}$ at the smallest length scales, given the turbulence description at the resolved scale. Thereby, there is no need for creation of a progress variable or new variables in the resolved scale. The analysis relates $\epsilon$ with the strain rate, vorticity, viscous dissipation rate, scalar gradients, scalar dissipation rate, and burning rate at the smallest turbulence length scales where diffusion-controlled burning is expected to be faster than at larger length scales and thereby dominant.
Two flamelet inflows with the same $\epsilon$ value at the resolved scale can still have statistical variations with a different pair of $S^{*}$ and $\omega$ values. So, the coefficients $C_{vd}$ and $C_{ke}$ have been used in the respective relations for $\epsilon$ with Kolmogorov-scale viscous dissipation rate and Kolmogorov-scale kinetic energy. Without further statistical refinement, the theory has the assumption of uniform coefficients over the resolved flow field. Conditions in the center of the flamelet can be identical, even for flamelets with different values for $\epsilon,S^{*},$ and $\omega$. For example, flamelets with the same values for maximum temperature $T_{max}$, scalar dissipation rate at the stoichiometric point in the flow $\chi_{st}$, and / or integrated burning rate can experience different resolved scale values such as $\epsilon$ or inflow values such as $S^{*}$ or $\omega$.
Direct numerical simulations in the literature [Nomura1992, Nomura1993, Boratav1996, Boratav1998, Betchov1956] indicate that values for Kolmogorov-scale vorticity and normal strain rate can be expected to have the same order of magnitude. With this consideration, we find that scalar gradients in the flamelet tend to scale with $\epsilon^{1/4}$. This means that scalar dissipation rate scales as $\epsilon^{1/2}$.